OpenStudy (anonymous):

hep!!!! fan and medal) solve: xy-4x+2y-8

OpenStudy (anonymous):

@ganeshie8 @mathstudent55

OpenStudy (baby_bear69):

Do you know how to combine like terms?

OpenStudy (anonymous):

no ;(

OpenStudy (jack1):

how do u mean... solve?

OpenStudy (jack1):

like factorise? find roots? find derivatives?

OpenStudy (anonymous):

@Baby_Bear69 are there?

OpenStudy (baby_bear69):

I'm trying to figure it out but my brain stopped working.. sorry .-.

OpenStudy (anonymous):

what are you being asked to solve for?

OpenStudy (anonymous):

it might help to know what subject are you studying

OpenStudy (anonymous):

look can ll me if I did right? xy-4x+2y-8 x(y-4)+2(y-4) (y-4)(x+2)

OpenStudy (anonymous):

if your being asked to factor then yes , thats right

OpenStudy (jack1):

FOIL first = x * y = xy = correct so far last = a * b = -8 sooo... outer and inner = ay + bx (x+a)(y+b) (x+2)(y-4) proof first = x * y = xy = correct so far last = a * b = 2 * -4 = -8 = correct outer and inner = ay + bx = 2y - 4x = correct

OpenStudy (anonymous):

thank you so much @jack @Jdosio

OpenStudy (anonymous):

np :D